OpenStudy (anonymous):

Fans & Medals!! I really need help!! The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –6 3 3 6 Line 2 x y –3 1 3 3 A. 0 B. exactly 1 C. exactly 2 D. infinitely many

OpenStudy (anonymous):

@ganeshie8 @dumbcow canyou guys help??

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@Ashleyisakitty

OpenStudy (anonymous):

@KendrickLamar2014 can you help?? plz

OpenStudy (dumbcow):

find slope of each line if slopes are different , then there will be only 1 solution if slopes are same, then check y-intercepts If same y-intercept, then lines are equal and there are infinite solutions If different, then no solution. the lines are parallel and never cross

OpenStudy (dumbcow):

Or just graph the lines and visually look to see how many times the lines cross :)

OpenStudy (anonymous):

yea my book was saying that can you show in example it doesnt have to be the same as the up^

OpenStudy (dumbcow):

you have to know the slope formula \[m = \frac{y_2 - y_1}{x_2 - x_1}\] for example: Line 1 has 2 points (-6,3) ---> (x1,y1) (3,6) ---->(x2, y2) \[m = \frac{6-3}{3- (-6)} = \frac{3}{9} = \frac{1}{3}\] Line 2: (-3,1) (3,3) \[m = \frac{3-1}{3-(-3)} = \frac{2}{6} = \frac{1}{3}\] They have same slope, find y-intercept use y=mx+b Line 1: (x,y) = (3,6) m = 1/3 \[6 = \frac{1}{3}(3) + b\] \[b = 6 - 1 = 5\] Line 2: (x,y) = (3,3) m=1/3 \[3 = \frac{1}{3}(3) + b\] \[b = 3-1 = 2\] Same slope, Different y-intercepts means No solution